After a 15% discount off of a pair of pants, the price is 14.50. What was the price before the discount?
Answer by april4869(19)

x=original price of the item, what you're solving for.
x(1-.15)=14.50.
it's 1-.15 because you don't want 15% of x, you want 15% OFF of x. so you really want 85% of x.
divide 14.50 by .85, and you'll have x.

(it's not 14.50 X 115% because then you're adding 15% of 14.50 to it, not 15% of the original amount you had. 15% of 14.50 does NOT equal 15% of x because x is larger.)
the pants are 17.06 NOT 16.68.
the way you can double check your answer is to take what you got for x and solve the problem backwards. so let's take $16.68 and show why it's wrong.
16.68*.15= 2.50200 (15% OFF 16.68)
16.68-2.50200=14.18 approx.
14.18 DOES NOT equal 14.50, which was the discounted price of the pants.
17.06*.15=2.559
17.06-2.559=14.50.
17.06 is the correct answer.